# Large Language Models Empowered Personalized Web Agents

### v1

@@ Appendix B Prompting Details
The prompt template for profile generation is shown in Figure 10 and Figure 11. And prompt template for instruction generation is shown in Figure 12, Figure 13, and Figure 14. Last, the prompt template for the user simulator is shown in Figure 15.
@@
User Profile Generation You will act as an online shopper. Given your time-series historical purchase information and corresponding reviews, you need to summarize and choose the most accurate and relevant option best describing you. During summarizing you should obey the following procedures: First, summarize your basic information, and choose or fill the most accurate and relevant option for each category. The categories and options are as follows: Gender: ¡GENDER¿. Age: ¡AGE¿. Occupation:¡OCCUPATION¿. Price Sensitivity: ¡PRICE SENSITIVITY¿. Shopping Interest: Summarize the product information. Brand Preference: Choose from the product information, and only keep brand names. Second, summarize your personal preferences across the following aspects: Diversity Preference: Do you prefer trying new things or sticking to familiar products? Choose from ¡DIVERSITY¿. Interaction Complexity Do you prefer simple and concise interactions, or do you enjoy detailed and thorough exchanges? Choose from ¡INTERACTION¿.. Tone and Style Summarize your overall emotional tone, speaking style, and expressive characteristics when giving reviews. Keep keywords. Item Reference Summarize your tendency to refer to specific products or brands in your reviews, like purchase history, shopping cart, or recommendations from friends. Keep keywords. Focus Aspect What aspects of products do you pay more attention to? Choose from ¡FOCUS ASPECT¿ or summarize others from reviews. Keep keywords. In the end, arrange all the above aspects using the JSON format, with each aspect as an individual key. Do not include any additional information or explanations and stay grounded. Your History: ¡HISTORY¿
@@
Details of Choices in Profile Generation ¡Gender¿: [Female, Male]. ¡AGE¿: [Under 18, 18-24, 25-34 , 35-44 , 45-49 , 50-55 , 56+]. ¡OCCUPATION¿: [Academic/Educator, Artist, Clerical/admin, College/grad student, Customer service, Doctor/health care, Executive/managerial, Farmer, Homemaker, K-12 student, Lawyer, Programmer, Retired, Sales/Marketing, Scientist, Self-employed, Technician/Engineer, Tradesman/Craftsman, Unemployed, Writer, Other]. ¡PRICE SENSITIVITY¿: [”High”: ”A Price-Conscious Shopper who is very sensitive to cost and seeks the best deals.”, ”Medium”: ”A Balanced Buyer who considers price but also values quality and features.”, ”Low”: ”A Value-Driven Consumer who prioritizes quality and features over price.”]. ¡DIVERSITY¿: [ ”High”: ”A Highly Adventurous Explorer eager to discover diverse products across categories. They often seek recommendations, and purchase a wide variety of items with varying ratings, and the user’s own ratings may often differ from the average. Their reviews are detailed and enthusiastic, reflecting their unique tastes and enjoyment of variety.”, ”Medium”: ”A Balanced Seeker who enjoys trying new products but also values familiarity. They appreciate targeted recommendations, purchase a moderate number of items with solid ratings and a reasonable number of ratings, and their reviews balance detailed feedback with concise, practical comments.”, ”Low”: ”A Meticulously Selective Buyer who sticks to tried-and-true products, showing little interest in new options. They purchase fewer items, favoring those with high ratings and a large number of ratings. Their own ratings are often very close to or slightly above the average, and their reviews are thoughtful and focused on familiar products.” ]. ¡INTERACTION¿: [ ”High”: ”A Thorough Conversationalist who enjoys detailed discussions, exploring all aspects of a product or service. They provide extensive reviews and value comprehensive support, engaging in multiple rounds of communication.”, ”Medium”: ”A Moderate Engager who balances simplicity with detail. They prefer clear communication but can engage in detailed exchanges when necessary. They provide reviews that are a mix of concise observations and some detailed insights, especially if they have strong feelings about a product.”, ”Low”: ”A Minimalist Interactor who values simplicity and efficiency. They prefer quick, straightforward interactions and leave brief, to-the-point reviews, focusing only on essential product aspects.” ]. ¡FOCUS ASPECT¿: [”Average Rating”, ”Number of Ratings”, ”Price”, ”Store”, ”Material”, ”Size”, ”Weight”, ”Brand”].
@@
Search Instruction Generation You will act as an online shopper. Your Profile: ¡PROFILE¿ You are looking for a product similar to the following product: ¡PRODUCT¿ You want to find a similar product, but you are not looking for an exact match. Generate a search request that is somewhat vague, reflecting your preferences and personalities without revealing the complete details of the target product. Rules: ¡DIVERSITY¿. ¡INTERACTION¿. You pay more attention to ¡FOCUS_ASPECT¿ of products, make sure to include some of them in the search request. Ensure the search request aligns with your overall tone and style: ¡TONE_AND_STYLE¿. Do not repeat the exact information in your profile or product. Instead, use your own words to convey the same information. Try to make the request as natural as possible and stick to the personalities in your profile. Do not include any additional information or explanations and stay grounded. Do not hallucinate information that is not provided, No more than ¡NUM¿ words.
@@
Recommendation Instruction Generation You will act as an online shopper. Your Profile: ¡PROFILE¿ You have recently shown interest in the following type of product: ¡PRODUCT¿ Now, you’re exploring options that could match your overall tastes, but you’re not sure exactly what you’re looking for. Generate a recommendation request that reflects your general preferences and style, but leaves room for flexibility and discovery. Rules: ¡DIVERSITY¿. ¡INTERACTION¿. You value ¡FOCUS_ASPECT¿ of products, but keep the request open-ended to allow for a variety of recommendations. Ensure the recommendation request aligns with your overall tone and style: ¡TONE_AND_STYLE¿. Do not restate your profile or the product. Use different words or hints to convey your preferences. Avoid being too specific or precise in your request. Try to make the request as natural as possible and stick to the personalities in your profile. Do not include any additional information or explanations and stay grounded. Do not hallucinate information that is not provided, No more than ¡NUM¿ words.
@@
Review Instruction Generation You will act as an online shopper. Your Profile: ¡PROFILE¿, You have recently purchased the following product: ¡PRODUCT¿, Your feelings about the product are: ¡REVIEW¿, Now, You want to write a review. Generate a review request to ask for assistance to create a complete review that reflects your preferences and typical review style. Rules: ¡INTERACTION¿. You value ¡FOCUS_ASPECT¿ of products, but keep the request open-ended to allow for a variety of recommendations. Ensure the review request aligns with your overall tone and style: ¡TONE_AND_STYLE¿. Do not simply restate your profile or feedback verbatim. Instead, paraphrase and expand to reflect a more comprehensive review. Try to make the request as natural as possible and stick to the personalities in your profile. Do not include any additional information or explanations and stay grounded. Do not hallucinate information that is not provided. No more than ¡NUM¿ words.
@@
User Simulation Instruction You are a user interacting with a personalized shopping agent. Your Profile: ¡PROFILE¿ You have purchased the following product: ¡PRODUCT¿ (In review tasks:) Your review is as follows: ¡REVIEW¿ The shopping agent will help you complete your shopping requests. Rules: Just generate one line at a time to simulate the user’s message. Do not hallucinate information that is not provided. Do not give additional instructions or ask questions, only respond to the agent’s questions. Do not provide any specific product details. Do not repeat the exact information in your profile or product. Instead, use your own words to convey the same information. Try to make the conversation as natural as possible and stick to the personalities in your profile. If the result is not satisfactory, you can express your dissatisfaction and provide clues to help the agent understand your preferences.

### v2

@@ Appendix C Prompting Details
The prompt template for profile generation is shown in Figure 10 and Figure 11. The prompt template for instruction generation is shown in Figure 12, Figure 13, and Figure 14. The prompt template for the user simulator is shown in Figure 15. Last, the prompt templates in task executing are shown in Figure 16 and Figure 17.
User Profile Generation
@@
You will act as an online shopper. Given your time-series historical purchase information and corresponding reviews, you need to summarize and choose the most accurate and relevant option best describing you. During summarizing you should obey the following procedures: First, summarize your basic information, and choose or fill the most accurate and relevant option for each category. The categories and options are as follows: Gender: ¡GENDER¿. Age: ¡AGE¿. Occupation:¡OCCUPATION¿. Price Sensitivity: ¡PRICE SENSITIVITY¿. Shopping Interest: Summarize the product information. Brand Preference: Choose from the product information, and only keep brand names. Second, summarize your personal preferences across the following aspects: Diversity Preference: Do you prefer trying new things or sticking to familiar products? Choose from ¡DIVERSITY¿. Interaction Complexity Do you prefer simple and concise interactions, or do you enjoy detailed and thorough exchanges? Choose from ¡INTERACTION¿. Tone and Style Summarize your overall emotional tone, speaking style, and expressive characteristics when giving reviews. Keep keywords. Item Reference Summarize your tendency to refer to specific products or brands in your reviews, like purchase history, shopping cart, or recommendations from friends. Keep keywords. Focus Aspect What aspects of products do you pay more attention to? Choose from ¡FOCUS ASPECT¿ or summarize others from reviews. Keep keywords. In the end, arrange all the above aspects using the JSON format, with each aspect as an individual key. Do not include any additional information or explanations and stay grounded. Your History: ¡HISTORY¿
@@
Details of Choices in Profile Generation
¡Gender¿: [Female, Male]. ¡AGE¿: [Under 18, 18-24, 25-34 , 35-44 , 45-49 , 50-55 , 56+]. ¡OCCUPATION¿: [Academic/Educator, Artist, Clerical/admin, College/grad student, Customer service, Doctor/health care, Executive/managerial, Farmer, Homemaker, K-12 student, Lawyer, Programmer, Retired, Sales/Marketing, Scientist, Self-employed, Technician/Engineer, Tradesman/Craftsman, Unemployed, Writer, Other]. ¡PRICE SENSITIVITY¿: [”High”: ”A Price-Conscious Shopper who is very sensitive to cost and seeks the best deals.”, ”Medium”: ”A Balanced Buyer who considers price but also values quality and features.”, ”Low”: ”A Value-Driven Consumer who prioritizes quality and features over price.”]. ¡DIVERSITY¿: [ ”High”: ”A Highly Adventurous Explorer eager to discover diverse products across categories. They often seek recommendations, and purchase a wide variety of items with varying ratings, and the user’s own ratings may often differ from the average. Their reviews are detailed and enthusiastic, reflecting their unique tastes and enjoyment of variety.”, ”Medium”: ”A Balanced Seeker who enjoys trying new products but also values familiarity. They appreciate targeted recommendations, purchase a moderate number of items with solid ratings and a reasonable number of ratings, and their reviews balance detailed feedback with concise, practical comments.”, ”Low”: ”A Meticulously Selective Buyer who sticks to tried-and-true products, showing little interest in new options. They purchase fewer items, favoring those with high ratings and a large number of ratings. Their own ratings are often very close to or slightly above the average, and their reviews are thoughtful and focused on familiar products.” ]. ¡INTERACTION¿: [ ”High”: ”A Thorough Conversationalist who enjoys detailed discussions, exploring all aspects of a product or service. They provide extensive reviews and value comprehensive support, engaging in multiple rounds of communication.”, ”Medium”: ”A Moderate Engager who balances simplicity with detail. They prefer clear communication but can engage in detailed exchanges when necessary. They provide reviews that are a mix of concise observations and some detailed insights, especially if they have strong feelings about a product.”, ”Low”: ”A Minimalist Interactor who values simplicity and efficiency. They prefer quick, straightforward interactions and leave brief, to-the-point reviews, focusing only on essential product aspects.” ]. ¡FOCUS ASPECT¿: [”Average Rating”, ”Number of Ratings”, ”Price”, ”Store”, ”Material”, ”Size”, ”Weight”, ”Brand”].
@@
Search Instruction Generation
You will act as an online shopper. Your Profile: ¡PROFILE¿ You are looking for a product similar to the following product: ¡PRODUCT¿ You want to find a similar product, but you are not looking for an exact match. Generate a search request that is somewhat vague, reflecting your preferences and personalities without revealing the complete details of the target product. Rules: ¡DIVERSITY¿. ¡INTERACTION¿. You pay more attention to ¡FOCUS_ASPECT¿ of products, make sure to include some of them in the search request. Ensure the search request aligns with your overall tone and style: ¡TONE_AND_STYLE¿. Do not repeat the exact information in your profile or product. Instead, use your own words to convey the same information. Try to make the request as natural as possible and stick to the personalities in your profile. Do not include any additional information or explanations and stay grounded. Do not hallucinate information that is not provided, No more than ¡NUM¿ words.
@@
Recommendation Instruction Generation
You will act as an online shopper. Your Profile: ¡PROFILE¿ You have recently shown interest in the following type of product: ¡PRODUCT¿ Now, you’re exploring options that could match your overall tastes, but you’re not sure exactly what you’re looking for. Generate a recommendation request that reflects your general preferences and style, but leaves room for flexibility and discovery. Rules: ¡DIVERSITY¿. ¡INTERACTION¿. You value ¡FOCUS_ASPECT¿ of products, but keep the request open-ended to allow for a variety of recommendations. Ensure the recommendation request aligns with your overall tone and style: ¡TONE_AND_STYLE¿. Do not restate your profile or the product. Use different words or hints to convey your preferences. Avoid being too specific or precise in your request. Try to make the request as natural as possible and stick to the personalities in your profile. Do not include any additional information or explanations and stay grounded. Do not hallucinate information that is not provided, No more than ¡NUM¿ words.
@@
Review Instruction Generation
You will act as an online shopper. Your Profile: ¡PROFILE¿, You have recently purchased the following product: ¡PRODUCT¿, Your feelings about the product are: ¡REVIEW¿, Now, You want to write a review. Generate a review request to ask for assistance to create a complete review that reflects your preferences and typical review style. Rules: ¡INTERACTION¿. You value ¡FOCUS_ASPECT¿ of products, but keep the request open-ended to allow for a variety of recommendations. Ensure the review request aligns with your overall tone and style: ¡TONE_AND_STYLE¿. Do not simply restate your profile or feedback verbatim. Instead, paraphrase and expand to reflect a more comprehensive review. Try to make the request as natural as possible and stick to the personalities in your profile. Do not include any additional information or explanations and stay grounded. Do not hallucinate information that is not provided. No more than ¡NUM¿ words.
@@
User Simulation Instruction
You are a user interacting with a personalized shopping agent. Your Profile: ¡PROFILE¿ You have purchased the following product: ¡PRODUCT¿ (In review tasks:) Your review is as follows: ¡REVIEW¿ The shopping agent will help you complete your shopping requests. Rules: Just generate one line at a time to simulate the user’s message. Do not hallucinate information that is not provided. Do not give additional instructions or ask questions, only respond to the agent’s questions. Do not provide any specific product details. Do not repeat the exact information in your profile or product. Instead, use your own words to convey the same information. Try to make the conversation as natural as possible and stick to the personalities in your profile. If the result is not satisfactory, you can express your dissatisfaction and provide clues to help the agent understand your preferences.
@@
Prompt in Single-Turn Track
As a personalized shopping agent, you can help users search for products, recommend products, or complete their reviews. Rules: The user will provide user_id and a request. You need to use the most appropriate tool to find the product or fill the review that matches the user’s request. You are not allowed to interact with the user. Make the best tool call based on the user’s request. You have only one chance to make a tool call, so make sure you have the best input for the tool. The tool will be provided, you need to use the tool and provide the most appropriate input for the tool. Do not use other tools. Formulate the best input for the tool based on the user’s request and the memory provided. Memory: (Depends on the method.) Functions to use: (Available functions and their descriptions.)
Figure 16. Prompt template for all methods in single-turn track.
Prompt in Multi-Turn Track
As a personalized shopping agent, you can help users search for products, recommend products, or complete their reviews. Rules: The user will provide user_id and a request. You need to use the tools to find the product or fill the review that matches the user’s request. The tool will be provided, you need to use the tool and provide the most appropriate input for the tool. Do not use any other tools. Formulate the best input for the tool based on the user’s request and the memory provided. You are allowed to interact with the user by ’respond’ to ask for more information or feedback, but steps are limited, and less steps are preferred. Your main goal is to help the user complete the task as accurately and efficiently as possible, do not keep responding to the user, focus on making the better tool calls. The evaluation will be based on the ranking of the target product in search and recommendation tasks, and the similarity of the review in the review task. When you think you have found the best input for the task tool calls, you can end the task by making a ’stop’ call. You should not make up any information or knowledge not provided from the user or the tools, or give subjective comments or recommendations. You should at most make one tool call at a time, and if you take a tool call, you should not respond to the user at the same time. If you respond to the user, you should not make a tool call. Memory: (Depends on the method.) Functions to use: (Available functions and their descriptions.)
Figure 17. Prompt template for all methods in multi-turn track.